# Neural Combinatorial Optimization Algorithms for Solving Vehicle Routing Problems: A Comprehensive Survey with Perspectives

### v2

@@ II-B2 Reinforcement Learning (RL)
Compared with SL-based solvers, RL ones autonomously learn to solve VRPs without human guidance and the effort of data labeling, thereby addressing the inherent limitation of SL’s reliance on high-quality labels. In addition, RL-based solvers consistently surpass SL-based solvers in terms of generalization due to RL’s interactive and exploratory nature [126]. Hence, following the study of [81], who introduced RL for training NNs to solve VRPs in 2017, a large number of RL-based solvers emerged [26, 27, 127, 79, 51, 128].
@@ III Learning to Construct (L2C) Solvers
Figure 3: Illustration of the generic construction process of L2C solvers, starting from an empty solution set and ending with complete solutions. Most L2C solvers are composed of an encoder and a decoder. Encoder is used to output the embeddings of VRP instances, while decoder selects nodes based on these embeddings to construct complete solutions.
@@ III-A Design of Encoders
In the field of NCO, Bello et al. [81] were the first to employ actor-critic RL for training Long Short-Term Memory (LSTM) [130] as an encoder, capturing the coordinate information of nodes and converting it into embeddings. Subsequently, to fully leverage the graph structure, Khalil et al. [82] proposed Graph Neural Network (GNN)-based S2V-DQN to embed node features. Differing from the previously employed RNN-based and GNN-based encoders, Kool et al. [26] proposed AM based on the superior Transformer architecture [131]. Specifically, Kool et al. adopted a total of three attention layers as an encoder, with each layer comprising a multi-head attention and a node-wise feed-forward sub-layer. Ever since, AM has become a prevalent encoder model choice, adopted by many L2C solvers [27, 79, 51, 132, 133, 134] as the backbone structure. For example, to pursue better performance, Kwon et al. [27] further extended the number of attention layers of AM from three to six. Furthermore, the following paragraphs present two primary Transformer-based encoder designs tailored for specific problem scales and VRPs.
@@
In summary, the superior performance of AM has established it as the predominant encoder type. While certain studies also consider the demands of specific VRP instances, thus tailoring the Transformer-based encoder design accordingly.
@@ III-B Design of Decoding Processes
The sequential nature of the AR approach imposes a fundamental speed limit on these L2C solvers [141]. As a result, these solvers struggle to construct high-quality solutions within short time constraints in real-world scenarios [7]. To expedite the decoding process, as shown in Fig 3, certain L2C solvers [5, 38] employ the Non-Autoregressive (NAR) approach, enabling the construction of complete solutions in a single decoding step. For example, Xiao et al. [38] employed the Greedy Search algorithm to decode the outputs of the encoder, i.e., the edge score matrix and the start node pointer, adhering to the order constraints inherent in TSP. Although NAR solvers excel in parallel solution constructions with low decoding latency, they often exhibit inferior performance in terms of quality. Subsequently, to obtain NAR solvers with low decoding latency and the capability to generate high-quality solutions, Xiao et al. [5] incorporated a knowledge distillation method to train a NAR solver (student model) using an AR solver (teacher model). For solvers utilizing NAR decoding methods, advocating for more sophisticated post-processing strategies, rather than solely relying on rudimentary Greedy Search algorithm, demonstrates a promising way of enhancing their performance.
@@ III-D Data Augmentation and Post-processing Strategies
In the field of Computer Vision and Natural Language Processing, data augmentation methods have been shown as instrumental in significantly elevating model performance [142, 143]. Similarly, data augmentation methods also play a pivotal role in the field of NCO, which can enrich the training datasets and enhance robustness [27, 35, 36, 128]. Specifically, to enrich the training datasets, Kwon et al. [27] were the first to introduce the symmetry coordinate transformation into L2C solvers as a data augmentation method. Subsequently, Kim et al. [79] proposed the orthogonal rotation coordinate transformation method to enrich the training datasets. Following [27, 79], Yao et al. [35] further extended the symmetry of VRPs, and proposed four coordinate transformation methods, namely rotation, symmetry, shrink, and noise. These data augmentation methods enrich the training datasets and remarkably contribute to the high-level performance of L2C solvers. This is particularly evident in the case of SL-based solvers, where a considerable quantity of training instances with high-quality labels is required.
@@
Concorde (2007) [144$\dagger$ 5.696 - 9m 7.765 - 43m - - - - - -
HGS (2022) [145$\dagger$ - - - - - - 10.366 - 1.2d 15.563 - 1.5d
@@
While L2C solvers can construct solutions within seconds using Greedy Search or Beam Search, they are highly prone to falling into the local optima [51]. Specifically, Greedy Search selects the node with the largest probability value at each decoding step, and Beam Search sequentially explores the top $k$ nodes based on their probabilities at each decoding step [25]. Therefore, other than data augmentation methods, post-processing strategies are also employed to pursue high-quality solutions with a sacrifice of computation time [81, 88, 90]. Initially, in the field of NCO, Bello et al. [81] proposed two kinds of post-processing strategies, namely Sampling and Active Search (AS). Specifically, Sampling randomly selects nodes utilizing the learned probability during the decoding process, constructs multiple candidate solutions, and ultimately reports the shortest among them. Adopting AS, the model actively updates its parameters for each test instance. Nevertheless, this process of adjusting the weight for each instance is both time-consuming and memory-intensive. To address this problem, Hottung et al. [88] further proposed Effective Active Search (EAS), which selectively adjusts a subset of model parameters. The adoption of EAS remarkably improves the performance of many L2C solvers (e.g., SCA [146]) without substantially increasing time and memory overhead. Subsequently, to couple with EAS while striving for SOTA performance, Choo et al. [56] proposed Simulation Guided Beam Search (SGBS). Specifically, in SGBS, rollouts are performed for the expanded child nodes of Beam Search to identify promising child nodes. Then, the child nodes with poor performance are pruned from consideration, and the beam is expanded to another depth.
@@ III-E Performance of L2C Solvers
Table III presents the performance of different L2C solvers on small-scale TSP and CVRP instances. Before 2019, most studies focused on enhancing the performance of L2C solvers from the perspective of model architecture. Then, with the establishment of AM [26] and POMO [27] in 2020, the success of these two studies had redirected subsequent research. Instead of exclusively focusing on improving encoders and decoders, there is a distinct shift towards advocating for data augmentation and post-processing strategies (e.g., SGBS [56]) to enhance the effectiveness of AM and POMO. Through the incorporation of advanced data augmentation and post-processing strategies, current L2I solvers still showcase performance on par with OR algorithms. For example, SGBS [56] attains gap values of 0.03% and 0.1% on the 100-node TSP and 100-node CVRP, respectively.However, existing L2C solvers have certain inadequacies when with faced specific situations. First, it is worth mentioning that certain L2C solvers prioritized high performance over solving time, leading to the conventional OR algorithms achieving shorter solution time [38]. This contradicts the inherent advantage of NCO solvers, i.e., their ability to solve VRP instances rapidly. Additionally, the generalization capability of these L2C solvers needs to be enhanced to better solve instances with different data distribution patterns and problem scales (see Section VI-A for more details). Finally, due to the time and memory complexity, existing Transformer-based L2C solvers struggle with handling large-scale VRP instances, e.g., the 10,000-node TSP (see Section VI-B for more details).
@@ IV-A Design of Regional Strategies
L2I solvers generally use a policy network to identify the regions that require repair (in another word, performance elevation) based on current solutions [6]. In the field of NCO, Chen et al. [39] first adopted a fully connected NN as a score predictor to identify a few nodes where rewriting might be beneficial. Subsequently, Costa et al. [41] and Sui et al. [147] adopted a network block consisting of GNN and RNN to select nodes based on current solutions. Considering the superior performance of Transformer-based networks for VRP, Wu et al. [40] proposed a Transformer-based network to choose node pairs for executing the 2-opt operator (see Section IV-B for more details about 2-opt operator). Subsequently, Ma et al. [78] proposed DACT to employ two Transformer networks to individually learn information from node coordinates and current solutions. Such information is then collaboratively used to determine the selection of node pairs. However, due to the adoption of two Transformer-based networks, the computing efficiency of DACT is not satisfactory [23]. To improve computing efficiency, Ma et al. [124] further proposed the Synthesis Attention mechanism, capable of simultaneously learning embeddings of node coordinates and current solutions to attain a comprehensive representation. While this regional strategy typically yields satisfactory results for small-scale VRPs, its efficiency diminishes when used to solve large-scale VRPs. This is because using NNs to select regions consisting of only a few nodes in large-scale VRPs could exponentially increase the overall solution time, ultimately leading to inefficient solution provision.
More recently, to better tackle large-scale VRP instances, certain L2I solvers [61, 148] propose a division strategy inspired by the D&C-based manner, rather than relying on NNs to select regions. Specifically, these D&C-based L2I solvers utilize the visiting order of the current complete solution to divide the problem into multiple sub-problems. Subsequently, all these sub-problems are simultaneously entered into the repair strategy for performance elevation. Finally, only the sub-problems showing improvement are considered for iteratively updating the current solutions. Meanwhile, to simultaneously improve multiple sub-problems, these L2I solvers [61, 148] also design a special repair strategy (see Section IV-B for more details). For these solvers, to prevent increased distance costs during the merging of sub-problems, establishing an efficient division strategy is crucial.
@@ IV-B Design of Repair Strategies
Currently, L2I solvers employ one of three distinct repair strategies to repair sub-tours of the selected regions. Initially, for simplicity, certain L2I solvers exploit the fixed operator throughout the entire iteration process to improve current solutions. For example, Chen et al. [39] adopted swap operator to iteratively improve current solutions, involving the exchange of the visiting order of two nodes, as shown in Fig. 5(a). Subsequently, Wu et al. [40] and Costa et al. [41] adopted 2-opt operator for iterative improvements in current solutions. As shown in Fig. 5(b), the 2-opt operator reconnects two new edges at a time. Despite the simplicity of swap and 2-opt operators, their performance falls short in searching for (sub-)optimal solutions within a limited timeframe (see Section IV-D).
To pursue better performance, certain L2I solvers exploit additional NNs to choose suitable operators for the regions selected by the regional strategy. For example, Lu et al. [93] exploited a Transformer-based network to select operators from two superior sets customized for CVRP, namely intra-route and inter-route sets. Specifically, intra-route operators endeavour to alter the order of visiting nodes within a single route, while inter-route ones attempt to exchange nodes among different routes. Meanwhile, Sui et al. [147] replaced 2-opt with 3-opt and proposed an NN to choose from seven different cases of the 3-opt operator, as depicted in Fig. 5(c). Generally speaking, the operator’s performance improves with a larger value of $k$ in $k$-opt operators, but simultaneously, the computing complexity significantly increases [42].
Finally, to tackle large-scale VRP instances, certain L2I solvers (e.g., SO [61] and GLOP [148]) exploit L2C solvers (e.g., AM [26]) to directly repair all sub-problems, bypassing operators like the 2-opt operator. The reason for employing L2C solvers as the repair strategy lies in the regional strategy utilized by these L2I solvers. These solvers divide large-scale problems into multiple smaller sub-problems, which can be efficiently solved by L2C solvers in parallel. In contrast, traditional operators lack the capability to simultaneously solve a batch of sub-problems. In addition, it is worth mentioning that while these solvers employ L2C solvers as their repair strategy, their key objective remains as iteratively improving the current complete solutions. Hence, we categorize these solvers into L2I categories.
@@
Methods TSP ($N=50$) TSP ($N=100$) CVRP ($N=50$) CVRP ($N=1,00$)
@@
Concorde (2007) [144$\dagger$ 5.696 - 9m 7.765 - 43m - - - - - -
HGS (2022) [145$\dagger$ - - - - - - 10.366 - 1.2d 15.563 - 1.5d
@@ IV-C5 Policy
The policy $\pi$ of the MDP for L2I solvers is parameterized by an NN in most cases. Unlike L2C solvers, the number of iterations of the L2I solvers $T$ could be any predefined value. Therefore, if $T$ is different, comparisons are unfair. Finally, the best solution identified across the entire iteration process is reported as the final solution.
@@ IV-D Performance of L2I Solvers
Table IV presents the performance of various L2I solvers on small-scale TSP and CVRP instances. Notably, NeuOpt [42], the most novel L2I solver utilizing the $k$-opt operator, outperforms the others, demonstrating only 0.02% and 0.6% gaps compared to Concorde and HGS, respectively. Specifically, to explicitly and automatically determine the values of $k$ for various cases, NeuOpt breaks down the sequential $k$-opt operator into a series comprising one starting move, several intermediate moves (determined by the proposed repair network), and one ending move. However, for L2I solvers, it is crucial to acknowledge that the problem of unfair comparisons becomes more pronounced when employing varying number of iterations $T$. For a fair comparison, adopting a fixed number of iterations for all L2I solvers is imperative [18]. In addition, the solution time also undergoes a significant rise with an increase of $T$. Finally, existing L2I solvers typically begin with randomly initiated feasible solutions and iteratively improve them for straightforwardness. However, this random initialization strategy hinders convergence and increases search overhead, particularly for large-scale VRP instances [149]. Future studies should explore more efficient strategies borrowed from the heuristic algorithms for constructing initial solutions to improve the time efficiency of L2I solvers.

### v3

@@ II-B2 Reinforcement Learning (RL)
Compared with SL-based solvers, RL ones autonomously learn to solve VRPs without human guidance and the effort of data labeling, thereby addressing the inherent limitation of SL’s reliance on high-quality labels. In addition, RL-based solvers consistently surpass SL-based solvers in terms of generalization due to RL’s interactive and exploratory nature [134]. Hence, following the study of [86], who introduced RL for training NNs to solve VRPs in 2017, a large number of RL-based solvers emerged [30, 17, 135, 84, 54, 136, 137, 138].
@@ III Learning to Construct (L2C) Solvers
Figure 3: Illustration of the generic construction process of L2C solvers, starting from an empty solution set and ending with complete solutions. Most L2C solvers are composed of an encoder and a decoder. The encoder is used to output the embeddings of VRP instances, while the decoder selects nodes based on these embeddings to construct complete solutions.
@@ III-A Design of Encoders
In the field of NCO, Bello et al. [86] were the first to employ actor-critic RL for training Long Short-Term Memory (LSTM) [140] as an encoder, capturing the coordinate information of nodes and converting it into embeddings. Subsequently, to fully leverage the graph structure, Khalil et al. [87] proposed Graph Neural Network (GNN)-based S2V-DQN to embed node features. Differing from the previously employed RNN-based and GNN-based encoders, Kool et al. [30] proposed AM based on the superior Transformer architecture [141]. Specifically, as shown in Fig. 4, Kool et al. adopted a total of three attention layers as an encoder, with each layer comprising a multi-head attention and a node-wise feed-forward sub-layer.
Figure 4: Illustration of AM [30].
Ever since, AM has become a prevalent encoder model choice, adopted by many L2C solvers [17, 84, 54, 142, 138, 143] as the backbone structure. For example, to pursue better performance, Kwon et al. [17] further extended the number of attention layers of AM from three to six. Furthermore, the following paragraphs present two primary Transformer-based encoder designs tailored for specific problem scales and VRPs.
@@
In summary, the superior performance of AM has established it as the predominant encoder type. While certain studies considered the demands of specific VRP instances, thus tailoring the Transformer-based encoder design accordingly.
@@ III-B Design of Decoding Processes
The sequential nature of the AR approach imposes a fundamental speed limit on these L2C solvers [150]. As a result, these solvers struggle to construct high-quality solutions within short time constraints in real-world scenarios [8]. To expedite the decoding process, as shown in Fig 3, certain L2C solvers [5, 41] employ the Non-Autoregressive (NAR) approach, enabling the construction of complete solutions in a single decoding step. For example, Xiao et al. [41] employed the Greedy Search algorithm to decode the outputs of the encoder, i.e., the edge score matrix and the start node pointer, adhering to the order constraints inherent in TSP. Although NAR solvers excel in parallel solution constructions with low decoding latency, they often exhibit inferior performance in terms of quality. Subsequently, to obtain NAR solvers with low decoding latency and the capability to generate high-quality solutions, Xiao et al. [5] incorporated a knowledge distillation method to train a NAR solver (student model) using an AR solver (teacher model). For solvers utilizing NAR decoding methods, advocating for more sophisticated post-processing strategies (see Section III-D for more details), rather than solely relying on rudimentary Greedy Search algorithm, demonstrates a promising way of enhancing their performance.
@@ III-D Data Augmentation and Post-processing Strategies
In the field of Computer Vision and Natural Language Processing, data augmentation methods have been shown as instrumental in significantly elevating model performance. Similarly, data augmentation methods also play a pivotal role in the field of NCO, which can enrich the training datasets and enhance robustness [17, 38, 39, 136]. Specifically, to enrich the training datasets, Kwon et al. [17] were the first to introduce the symmetry coordinate transformation into L2C solvers as a data augmentation method. Subsequently, Kim et al. [84] proposed the orthogonal rotation coordinate transformation method to enrich the training datasets. Following [17, 84], Yao et al. [38] further extended the symmetry of VRPs, and proposed four coordinate transformation methods, namely rotation, symmetry, shrink, and noise. These data augmentation methods enrich the training datasets and remarkably contribute to the high-level performance of L2C solvers. This is particularly evident in the case of SL-based solvers, where a considerable quantity of training instances with high-quality labels is required.
@@
Concorde (2007) [151$\dagger$ 5.696 - 9m 7.765 - 43m - - - - - -
HGS (2022) [152$\dagger$ - - - - - - 10.366 - 1.2d 15.563 - 1.5d
@@
While L2C solvers can construct solutions within seconds using Greedy Search or Beam Search, they are highly prone to falling into the local optima [54]. Specifically, Greedy Search selects the node with the largest probability value at each decoding step, and Beam Search sequentially explores the top $k$ nodes based on their probabilities at each decoding step [29]. Therefore, other than data augmentation methods, post-processing strategies are also employed to pursue high-quality solutions with a sacrifice of computation time [86, 93, 95, 154]. Initially, in the field of NCO, Bello et al. [86] proposed two kinds of post-processing strategies, namely Sampling and Active Search (AS). Specifically, Sampling randomly selects nodes utilizing the learned probability during the decoding process, constructs multiple candidate solutions, and ultimately reports the shortest among them. Adopting AS, the model actively updates its parameters for each test instance. Nevertheless, this process of adjusting the weight for each instance is both time-consuming and memory-intensive. To address this problem, Hottung et al. [93] further proposed Effective Active Search (EAS), which selectively adjusts a subset of model parameters. The adoption of EAS remarkably improves the performance of many L2C solvers (e.g., SCA [155]) without substantially increasing time and memory overhead. Subsequently, to couple with EAS while striving for SOTA performance, Choo et al. [153] proposed Simulation Guided Beam Search (SGBS). Specifically, in SGBS, rollouts are performed for the expanded child nodes of Beam Search to identify promising child nodes. Then, the child nodes with poor performance are pruned from consideration, and the beam is expanded to another depth. Recently, Luo et al. [39] proposed the Random Re-Construct (RRC) post-processing strategy, which resembles the iterative ruin-and-repair approach used in heuristic algorithms. Specifically, RRC randomly samples a partial solution from the constructed complete solution and restructures it using the proposed NCO solver to obtain a new partial solution.
@@ III-E Performance of L2C Solvers
Table III presents the performance of different L2C solvers on small-scale TSP and CVRP instances. Before 2019, most studies focused on enhancing the performance of L2C solvers from the perspective of model architecture. Then, with the establishment of AM [30] and POMO [17] in 2020, the success of these two studies has redirected subsequent research. Instead of exclusively focusing on improving encoders and decoders, there is a distinct shift towards advocating for data augmentation and post-processing strategies (e.g., SGBS [153]) to enhance the effectiveness of AM and POMO. Through the incorporation of advanced data augmentation and post-processing strategies, current L2C solvers still showcase performance on par with OR algorithms. For example, SGBS [153] attains gap values of 0.03% and 0.1% on the 100-node TSP and 100-node CVRP, respectively. However, existing L2C solvers have certain inadequacies when with faced specific situations. First, it is worth mentioning that certain L2C solvers prioritized high performance over solving time, leading to the conventional OR algorithms achieving shorter solution time [41]. This contradicts the inherent advantage of NCO solvers, i.e., their ability to solve VRP instances rapidly. Additionally, the generalization capability of these L2C solvers needs to be enhanced to better solve instances with different data distribution patterns and problem scales (see Section VI-A for more details). Finally, due to the time and memory complexity, existing Transformer-based L2C solvers struggle with handling large-scale VRP instances, e.g., the 10,000-node TSP (see Section VI-B for more details).
@@ IV-A Design of Regional Strategies
L2I solvers generally use a policy network to identify the regions that require repair (in another word, performance elevation) based on current solutions [6]. In the field of NCO, Chen and Tian [42] first adopted a fully connected NN as a score predictor to identify a few nodes where rewriting might be beneficial. Subsequently, Costa et al. [44] and Sui et al. [156] adopted a network block consisting of GNN and RNN to select nodes based on current solutions. Considering the superior performance of Transformer-based networks for VRP, Wu et al. [43] proposed a Transformer-based network to choose node pairs for executing the 2-opt operator (see Section IV-B for more details about 2-opt operator). Subsequently, Ma et al. [83] proposed DACT to employ two Transformer networks to individually learn information from node coordinates and current solutions. Such information is then collaboratively used to determine the selection of node pairs. However, due to the adoption of two Transformer-based networks, the computing efficiency of DACT is not satisfactory [27]. To improve computing efficiency, Ma et al. [132] further proposed the Synthesis Attention mechanism, capable of simultaneously learning embeddings of node coordinates and current solutions to attain a comprehensive representation. While this regional strategy typically yields satisfactory results for small-scale VRPs, its efficiency diminishes when used to solve large-scale VRPs. This is because using NNs to select regions consisting of only a few nodes in large-scale VRPs could exponentially increase the overall solution time, ultimately leading to inefficient solution provision.
To better tackle large-scale VRP instances, certain L2I solvers [157] propose a division strategy inspired by the D&C-based manner, rather than relying on NNs to select regions. Specifically, these D&C-based L2I solvers utilize the visiting order of the current complete solution to divide the problem into multiple sub-problems. Subsequently, all these sub-problems are simultaneously entered into the repair strategy for performance elevation. Finally, only the sub-problems showing improvement are considered for iteratively updating the current solutions. Moreover, to simultaneously improve multiple sub-problems, these L2I solvers [64, 157] also design a special repair strategy (see Section IV-B for more details). For these solvers, to prevent increased distance costs during the merging of sub-problems, establishing an efficient division strategy is crucial.
@@ IV-B Design of Repair Strategies
Currently, L2I solvers employ one of three distinct repair strategies to repair sub-tours of the selected regions. Initially, for simplicity, certain L2I solvers exploit the fixed operator throughout the entire iteration process to improve current solutions. For example, Chen and Tian [42] adopted swap operator to iteratively improve current solutions, involving the exchange of the visiting order of two nodes, as shown in Fig. 6(a). Subsequently, Wu et al. [43] and Costa et al. [44] adopted 2-opt operator for iterative improvements in current solutions. As shown in Fig. 6(b), the 2-opt operator reconnects two new edges at a time. Despite the simplicity of swap and 2-opt operators, their performance falls short in searching for (sub-)optimal solutions within a limited timeframe (see Section IV-D).
To pursue better performance, certain L2I solvers exploit additional NNs to choose suitable operators for the regions selected by the regional strategy. For example, Lu et al. [98] exploited a Transformer-based network to select operators from two superior sets customized for CVRP, namely intra-route and inter-route sets. Specifically, intra-route operators endeavour to alter the order of visiting nodes within a single route, while inter-route ones attempt to exchange nodes among different routes. Meanwhile, Sui et al. [156] replaced 2-opt with 3-opt and proposed an NN to choose from seven different cases of the 3-opt operator, as depicted in Fig. 6(c). Generally speaking, the operator’s performance improves with a larger value of $k$ in $k$-opt operators, but simultaneously, the computing complexity significantly increases. To automatically determine the values of $k$ at different iteration stages, Ma et al. [45] proposed NeuOpt, which decomposes the sequential $k$-opt operator into a structured sequence consisting of a starting move, $k$ intermediate moves (determined by the proposed repair network), and an ending move. This design effectively balances coarse-grained (larger $k$) and fine-grained (smaller $k$) search strategies.
Finally, to tackle large-scale VRP instances, certain L2I solvers (e.g., SO [64] and GLOP [157]) exploit L2C solvers (e.g., AM [30]) to directly repair all sub-problems, bypassing operators like the 2-opt operator. The reason for employing L2C solvers as the repair strategy lies in the regional strategy utilized by these L2I solvers. These solvers divide large-scale problems into multiple smaller sub-problems, which can be efficiently solved by L2C solvers in parallel. In contrast, traditional operators lack the capability to simultaneously solve a batch of sub-problems. In addition, Li et al. [158] proposed DRHG, which first decomposes the complete solution into certain isolated nodes and segments (with each segment composed of multiple consecutive nodes). Subsequently, DRHG simplifies the segments into hyperedges (temporally retaining only the two endpoints) and utilizes the L2C solver to connect the endpoints of these hyperedges with the isolated points, ultimately reconstructing a complete solution. This design directs the L2C solver’s attention to the disrupted regions while reducing the encoding complexity of all nodes within the segments, particularly suited to handle large-scale VRP instances. In addition, it is worth mentioning that while these solvers employ L2C solvers as their repair strategy, their key objective remains as iteratively improving the current complete solutions. Hence, we categorize these solvers into L2I categories.
@@
Methods TSP ($N=50$) TSP ($N=100$) CVRP ($N=50$) CVRP ($N=100$)
@@
Concorde (2007) [151$\dagger$ 5.696 - 9m 7.765 - 43m - - - - - -
HGS (2022) [152$\dagger$ - - - - - - 10.366 - 1.2d 15.563 - 1.5d
@@ IV-C5 Policy
The policy $\pi$ of the MDP for L2I solvers is parameterized by an NN in most cases. Unlike L2C solvers, the number of iterations of the L2I solvers $T$ could be any predefined value. Therefore, if $T$ varies, comparing different L2I solvers becomes inequitable. Finally, the best solution identified across the entire iteration process is reported as the final solution.
@@ IV-D Performance of L2I Solvers
Table IV presents the performance of various L2I solvers on small-scale TSP and CVRP instances. Notably, NeuOpt [45], the most novel L2I solver utilizing the $k$-opt operator, outperforms the others, demonstrating only 0.02% and 0.6% gaps compared to Concorde and HGS, respectively. However, for L2I solvers, it is crucial to acknowledge that the problem of unfair comparisons becomes more pronounced when employing varying number of iterations $T$. For a fair comparison, adopting a fixed number of iterations for all L2I solvers is imperative [21]. In addition, the solution time also undergoes a significant rise with an increase of $T$. Finally, existing L2I solvers typically begin with randomly initiated feasible solutions and iteratively improve them for straightforwardness. However, this random initialization strategy hinders convergence and increases search overhead, particularly for large-scale VRP instances [159]. Future studies should explore more efficient strategies borrowed from the heuristic algorithms for constructing initial solutions to improve the time efficiency of L2I solvers.